OpenStudy (anonymous):

Select all that apply. If 2x²-6x=0, then

OpenStudy (amorfide):

take out everything 2x^2 and 6x have in common so 2 goes into 2 and 6 2(x^2 - 3x)=0 they both have an x 2x(x-3)=0 now you have two equations for x 2x=0 or x-3=0

OpenStudy (anonymous):

ok thanks a lot